OpenStudy (anonymous):

find the y intercept and x intercept for the equation 5x+2y=30

hero (hero):

1. Set y = 0, then solve for x 2. Set x = 0, then solve for y

OpenStudy (anonymous):

5x - 2y = 30 5(0) - 2y = 30 y = -15 (y-intercept) Take the equation, set y = 0, and solve for x to get the x-intercept 5x - 2(0) = 30 x = 6 (x-intercept)

OpenStudy (anonymous):

y=0, x=6 x=0, y=-15

OpenStudy (anonymous):

X and Y intercepts are when the x or y counterpart is equal to zero. So: for the x-intercept: 2(0)=30-5x for the y-intercept 5(0)=30-2y If you need more clarification feel free to ask!

OpenStudy (anonymous):

5x+2y=30 y-intercept x=0 5(0)+2y=30 2y=30 y=15 x-intercept y=0 5x+2(0)=30 x=6